OpenStudy (anonymous):

MEDAL AND FAN PLEASE HELP Cylinder A has a radius of 7 inches and a height of 5 inches. Cylinder B has a volume of 490π. What is the percent change in volume between cylinders A and B? (5 points)

OpenStudy (anonymous):

@Preetha @pooja195

OpenStudy (mindblast3r):

all you need to to is find the both of the cylinder volumes.

OpenStudy (mindblast3r):

then divide the first cylinder by the second one, that will give you a number, you then turn that number into a percentage.

OpenStudy (anonymous):

im totally lost here can you tell me how to find the volumes and stuff

OpenStudy (mindblast3r):

ok

OpenStudy (anonymous):

Cylinder B is 50% smaller than cylinder A. Cylinder B is 75% bigger than cylinder A. Cylinder B is 150% smaller than cylinder A. Cylinder B is 100% bigger than cylinder A.

OpenStudy (mindblast3r):

\[V=\pi r^2h\]

OpenStudy (mathstudent55):

Since you were given the volume of cylinder B in terms of pi, calculate the volume of cylinder A also in terms of pi.

OpenStudy (mindblast3r):

\[V=\pi(7)^2(5)=245inches^3\]

OpenStudy (mindblast3r):

\[Cylinder B = 490\pi\]

OpenStudy (anonymous):

so what would be my final answer? @Mindblast3r

OpenStudy (mathstudent55):

\(\Large V_B = 490 \pi~in.^3\) \(\Large V_A = \pi r^2 h = \pi (7~in.)^2(5 ~in.) = 245 \pi ~in.^3\)

OpenStudy (anonymous):

what do i do after? @mathstudent55

OpenStudy (mathstudent55):

Since both volumes are shown in terms of pi, and you are interested in the percent difference, you can drop pi from both. What is the percent change from 245 to 490?

OpenStudy (mindblast3r):

\(\color{blue}{\text{Originally Posted by}}\) @Mindblast3r \[\frac{ CylinderA }{ CylinderB }=\frac{ 245\pi }{ 490\pi }=0.5=50percent\] \(\color{blue}{\text{End of Quote}}\)

OpenStudy (anonymous):

Thank you!!! @Mindblast3r

OpenStudy (anonymous):

I got it thank you both @mathstudent55 @Mindblast3r

OpenStudy (mathstudent55):

@Mindblast3r You need to read the question carefully and make sure you answer the question.

OpenStudy (mindblast3r):

wait i think

OpenStudy (anonymous):

It would be A right?

OpenStudy (mathstudent55):

The answer is not 50%

OpenStudy (mindblast3r):

oops no wait.

OpenStudy (anonymous):

oh lol never mind

OpenStudy (mathstudent55):

\(percent~change = \dfrac{new~amount - old~amount}{old~amount} \times 100\)

OpenStudy (anonymous):

Which is it theN? @mathstudent55

OpenStudy (mathstudent55):

Use the formula above to find a percent change. Use 245 as the old amount and 490 as the new amount.

OpenStudy (anonymous):

Its D right????? @mathstudent55

OpenStudy (anonymous):

100 percent

OpenStudy (mathstudent55):

\(percent~change = \dfrac{new~amount - old~amount}{old~amount} \times 100\) \(percent~change = \dfrac{490 - 245}{245} \times 100\) \(percent~change = \dfrac{245}{245} \times 100\) \(percent~change = 1 \times 100\) percent change = 100%

OpenStudy (anonymous):

THANK YOUUUU

OpenStudy (mathstudent55):

Correct. The answer is that cylinder B is 100% bigger than cylinder A.

OpenStudy (anonymous):

Can you help me with another one? please @mathstudent55

OpenStudy (mathstudent55):

@Mindblast3r Percent change is based on the formula I wrote above. Percent comparisons are tricky. If the question were "what percent of cylinder B's volume is cylinder A?" Then the answer would be 50%. The way the question was asked, though, it requires an answer of 100%. This is why you need to read the problem carefully bec with he given info, the answer could be 50% or 100% or other numbers depending on which way the question is asked.

OpenStudy (mathstudent55):

Sorry, it very late for me. GTG

OpenStudy (mindblast3r):

ahh i see what they did.